# Language Models are Crossword Solvers

### v2

@@ Abstract
Crosswords are a form of word puzzle that require a solver to demonstrate a high degree of proficiency in natural language understanding, wordplay, reasoning, and world knowledge, along with adherence to character and length constraints. In this paper we tackle the challenge of solving crosswords with Large Language Models (LLMs). We demonstrate that the current generation of state-of-the art (SoTA) language models show significant competence at deciphering cryptic crossword clues, and outperform previously reported SoTA results by a factor of 2-3 in relevant benchmarks. We also develop a search algorithm that builds off this performance to tackle the problem of solving full crossword grids with LLMs for the very first time, achieving an accuracy of 93% on New York Times crossword puzzles. Contrary to previous work in this area which concluded that LLMs lag human expert performance significantly, our research suggests this gap is a lot narrower.
@@
Soumadeep Saha  and Sutanoya Chakraborty  and Saptarshi Saha  and Utpal Garain Indian Statistical Institute Kolkata, India Correspondence: soumadeep.saha97@gmail.com
@@ 1 Introduction
Crossword puzzles are a type of word game that typically takes the form of a square grid of white and black squares. The objective of the puzzle is to fill the white squares with letters from words or phrases based on the provided clues (see Figure 1). Although crosswords come in a variety of styles, the two most popular ones are the American style or straight crosswords and cryptic crosswords.
@@
Figure 1: Example of a crossword puzzle (left) and cryptic clues (right). (left) The grid must be filled up with answers from the semantic clues provided. The gray highlighted squares produce additional constraints, e.g. first character of answer to clue 1 (across) must be the same as the first character of answer to clue 4 (down). Example by Fred Piscop. (right) In cryptic crosswords, the clues involve some form of wordplay and synonyms, and often involve world knowledge. Examples are taken from the cryptonite dataset.
Modern LLMs demonstrate astounding skills in reasoning, coding, wordplay, question answering (QA), and a multitude of other tasks (Wei et al., 2022a; Minaee et al., 2024). In this paper we attempt to analyze LLMs’ ability to solve crossword puzzles, with the primary goal of understanding strengths and weaknesses demonstrated by SoTA LLMs.
Solving crossword puzzles requires high levels of proficiency in understanding contextual clues, wordplay, world knowledge, and reasoning (see Figure 1 (right)). To illustrate the kinds of skills involved, consider as an example the cryptic crossword clue - ‘‘Culminating point of story about Judy’s husband by the railway track (5,4)’’.  [^1] Example taken from https://lovattspuzzles.com. Solving this clue requires connecting ‘Judy’ to the popular puppet show Punch and Judy, and inferring that ‘Judy’s husband’ refers to PUNCH. Additionally, we must observe that ‘railway track’ is synonymous to LINE, and combining these gives PUNCHLINE which also means ‘Culminating point of story’.
There is another aspect to solving crossword puzzles, which is constraint satisfaction (see Figure 1). Not only does the generated output have to adhere to the length limitations, it also has to incorporate character level constraints from previously generated answers. Successful crossword solvers generate answers based on clues, length information, character information from earlier answers, etc. and have the ability to backtrack when faced with inconsistent solutions in a targeted manner. Thus, even when the crossword clues are simpler, as in straight crosswords, the problem still poses a significant challenge to SoTA LLMs. Since solving crosswords requires proficiency in several desirable areas, our objective with this paper is to probe LLMs’ abilities at this multi-faceted task.
@@
To that end, we first perform extensive analysis to understand how well LLMs can solve crossword clues based on the provided contextual information and constraints. We show that current-generation SoTA LLMs demonstrate massively improved performance when compared to previous SoTA baselines, often outperforming them by a factor of 2 or 3, without any fine-tuning.
@@
In our analysis, we found that the ability of LLMs to adhere to length constraints, which is a critical component of solving crossword puzzles, is somewhat limited. We present results demonstrating that an LLM’s ability to count the number of constituent characters in a word degrades with the prevalence of the word. This suggests that their ability to perform this task does not generalize well.
@@
Armed with the knowledge gathered from our analysis, we tackle the full problem - i.e. not only answering clues, but filling in the grid considering all the imposed constraints. We device an algorithm called SweepClip which can solve straight crossword puzzles with an accuracy of 93.1% with the aid of an LLM. This algorithm exploits the constraints imposed by previously generated answers to improve future answers, and achieves results comparable to rule-based systems. This algorithm also improves the clue deciphering capability of LLMs, more than doubling the initial performance.
@@
To check if the performance improvements shown by LLMs is due to dataset contamination, we perform tests with post knowledge-cutoff benchmarks and human evaluation. We believe our work has been a major step towards solving crosswords with LLMs, and with future advances in certain key areas, LLMs’ performance in this task will be comparable to, if not better than human experts.
@@ 2 Background
Traditional approaches to solving crossword puzzles typically involve two parts - a candidate answer proposal system and a grid-filling algorithm (Angelini et al., 2005; Ginsberg, 2011). One such system, Proverb, Littman et al. (2002) achieves a letter accuracy of $98.1\%$ on crosswords from the New York Times (NYT). The candidate answer proposal systems rely on large datasets of crossword clues and answers (Shazeer et al., 1999) and grid-filling relies on algorithms for the constraint satisfaction problem (Angelini et al., 2005; Zugarini et al., 2023; Shazeer et al., 1999; Kulshreshtha et al., 2022). Kulshreshtha et al. (2022) who curated a dataset of NYT crossword puzzles and established benchmarks with language models have highlighted this task as “… a new high bar for AI systems”.
@@
Attempts to solve cryptic crossword clues with traditional algorithms, employing a large dataset of clues and a curated context free grammar (CFG) based parser (Deits, 2015), were not found to be performant ($7\%$ accuracy according to Sadallah et al. (2024)). Using LLMs to solve cryptic crossword clues has received attention in some recent papers. Efrat et al. (2021) curated a large dataset of cryptic crossword clues sourced from two UK dailies - The Times and The Telegraph and fine-tuned a T5-Large (Raffel et al., 2020) model to establish baseline performances. They proposed a training split for their dataset where the answers in the training and test sets are mutually exclusive so the model cannot rely on memorization. Following this, we have the work by Rozner et al. (2021), where they curated a dataset of cryptic crossword clues from The Guardian, and fine-tuned a T5-Large model with a curriculum learning approach, demonstrating performance improvements. Rozner et al. (2021) had some salient criticisms of the work by Efrat et al. (2021), and suggested that a disjoint train-test split is not a strong enough condition to ensure that the models learn to solve cryptic crosswords, since the models demonstrate “…robustness to plural and other inflections”. Rozner et al. (2021) proposed a split where words with similar roots are grouped together, and demonstrated that this stringent criteria leads to reduced performance.
@@
The most recent work by Sadallah et al. (2024) presents results with recent LLMs like Mistral-7B (Jiang et al., 2023), LLaMA2-7B (Touvron et al., 2023) and ChatGPT  [^2] It is unclear to us which version is used.(OpenAI, 2021) in few-shot settings and also by fine-tuning the Mistral model. They report that ChatGPT surpasses other models with an accuracy of $9.5\%$, which demonstrates a significant performance gap with respect to human experts, who solve $99\%$ of cryptic crossword clues (Sadallah et al., 2024). They point out key limitations in their work, like the extremely limited set of LLMs used, and potential for data contamination.
All three recent works on cryptic crosswords conclude that there is a massive gap in performance between LLMs and human experts. Also, these recent works focusing on solving cryptic crosswords with LLMs do not attempt to solve the full problem, i.e. they do not attempt to incoporate grid information and the constraints imposed by them, and rather elect to pose the problem of solving crosswords in the form of a question answering (QA) task. Kulshreshtha et al. (2022) attempted to solve NYT crossword puzzles with language models and a Satisfiability Modulo Theories (SMT) solver with limited success, and had to resort to culling the crossword grid based on candidate generations and ground truth answers.
In this paper, we undertake the challenge of solving a full crossword puzzle by developing an algorithm that incorporates grid information based on the answers given by an LLM. We find that with the application of our algorithm, the base performance of LLMs can be boosted significantly, and that the gap between the performance of LLMs and human experts is much smaller than previously anticipated.
@@ 3 Analyzing Crossword Clue Solving
In this section we analyze the performance of several LLMs at different scales at the clue deciphering task.
@@ 3.1 Datasets
Most of our analysis is performed on three crossword puzzle datasets. The first one covering straight or American style crosswords was curated by us from the very popular and long running puzzle section of the NYT. The other two, namely, Cryptonite by Efrat et al. (2021) and word-init-disjoint by Rozner et al. (2021) cover cryptic crossword puzzles. As discussed in Section 2, the two cryptic crossword datasets have methodological differences in how they partition training-test splits, however, this is not a very pertinent consideration for us, since we do not perform any training. A bulk of the results are reported on the NYT dataset we curated for straight crosswords and word-init-disjoint (abbreviated as Init) for cryptic crosswords since it was found to be more challenging than Cryptonite for LLMs. See Appendix B for more details.
@@ 3.2 Answering Crossword Clues
The first step in solving a crossword puzzle is deciphering its clues, so following in the scheme of Rozner et al. (2021); Sadallah et al. (2024); Kulshreshtha et al. (2022) and Efrat et al. (2021) we first explore this as a QA task, where the clue and its length is provided to the LLM as a query, with the expectation that it will produce the corresponding answer.
We report results with Phi 3 3.8B Instruct (Abdin et al., 2024), Mistral 7B Instruct (Jiang et al., 2023), LlaMa 2 70B (Touvron et al., 2023), LlaMa 3 8B Instruct, Llama 3 70B (Meta, 2024), Mixtral 8x7B (Jiang et al., 2024), Claude 3 Sonnet (Anthropic, 2024), GPT 3.5 Turbo, and GPT 4 Turbo (OpenAI, 2023) to cover a wide range of parameter scales and a mix of open-weights and proprietary models. We investigate the performance of these models with few-shot prompts (5-shot, 10-shot) on samples from the NYT dataset (ours), Cryptonite and the word-init-disjoint split (see Figure 2).
Figure 2: Analyzing LLMs’ ability to generate answers from crossword clues. We test LLMs at different scales on the NYT, Cryptonite and word-init-disjoint dataset with 5-shot, 10-shot prompts. We show that current generation of LLMs show significantly improved performance, and in particular GPT 4 Turbo outperforms the previous SoTA by 2$\sim$ 2$\times$ 2 × on the word-init-disjoint dataset with 5-shot prompts.
We find that the performance difference between 5-shot and 10-shot  [^3] We elected not to perform experiments with 10-shot prompts on Claude and GPT-4-Turbo due to budget constraints. prompted answers is not significant (see Figure 2), which is consistent with the findings by Sadallah et al. (2024). We also note that there is an appreciable difference (up to 5%) in performance of models across the board between Cryptonite and Init, with LLMs showing diminished performance on the Init dataset.
The models demonstrate improved performance with scale across datasets, and in particular show impressive performance on the NYT dataset with Llama 3 70B, GPT 3.5 Turbo, Claude 3 Sonnet, and GPT-4-Turbo achieving 27.2%, 26.05%, 37.7% and 41.2% accuracy (exact match) respectively.
The performance of LLMs on cryptic crosswords is worse compared to straight crosswords, however, Claude 3 Sonnet and GPT-4-Turbo outperform previous SoTA results on both the cryptic crossword datasets, achieving an accuracy of 12.9%, 23.5% on Cryptonite and 10.8%, 18.7% on Init respectively. The performance of GPT-4-Turbo is rather extraordinary, with a 1.97$\times$ improvement over previous SoTA (9.5% reported by Sadallah et al. (2024), see Figure 2). LlaMa 3 70B and GPT 3.5 Turbo demonstrate performance comparable to rule-based systems, and the smaller LLMs we tested perform far worse.
Further details about experimental protocols can be found in Appendix C and the details of the results summarized in Figure 2 can be found in Appendix F.
@@ 3.3 Exploiting Partially Filled Grids
In the course of solving a crossword puzzle, we encounter intermediate states where some of the clues have been deciphered. Successful crossword solvers exploit these clues in order to inform their decisions about future answers. For example, in Figure 1 (left), when we want to solve for the clue in position 14 (across), we can use the characters from position 2 (down) and 4 (down) to narrow down the set of possible answers to only those that fit the template “_ T _ P”. In this section we study LLMs’ performance at exploiting these constraints to arrive at better answers (see Table 1).
For this experiment we chose the best performing open-weights and proprietary model i.e. LLaMA 3 70B and GPT-4-Turbo respectively. We also report results on some smaller models to see if the trends hold. We report results on the NYT dataset and Init dataset with 5-shot prompts. For each query, $k$of the characters (letters) of the answer is provided alongside the clue and expected length of the answer, and the LLMs are expected to “unmask” the remaining characters using the provided constraint information and the crossword clue.  [^4] Further experimental details can be found in Appendix C.
@@
Mistral 7B (fine-tuned) 27.0%
Table 1: LLMs can exploit information provided by solving other clues to improve their accuracy. During crossword solving, we encounter partially filled grids, which provides hints about potential solutions in the form of character constraints. We test LLMs to see if they can incorporate this information and demonstrate improved performance. $\dagger$ is the results reported by Sadallah et al. (2024) by fine-tuning a Mistral 7B model on cryptic crossword solving. We demonstrate that models can indeed exploit these constraints and we find that GPT-4-Turbo with 5-shot prompts outperforms the fine-tuned Mistral 7B model by a factor of $\sim$2.8$\times$.
We observe (Table 1) that, in all but one case, LLMs show improved performance with an increasing percentage of constraint information for both datasets. Additionally, to compare the performance of GPT-4-Turbo to previously reported SoTA results by Sadallah et al. (2024) we perform the experiment with the same settings and dataset split as them, to find that GPT-4-Turbo with 5 shot prompts (76.3% accuracy) outperforms the fine-tuned Mistral 7B model (27% accuracy) by a factor of 2.8$\times$.
The fact that LLMs can successfully exploit constraints to answer crossword clues better, suggests that they are well suited to the task of solving full crosswords. The massive jump in performance observed in modern SoTA LLMs at this task and the task in Section 3.2 is extremely serendipitous, and we investigate this in further detail in Section 5.3.

### v3

@@ Abstract
Crosswords are a form of word puzzle that require a solver to demonstrate a high degree of proficiency in natural language understanding, wordplay, reasoning, and world knowledge, along with adherence to character and length constraints. In this paper we tackle the challenge of solving crosswords with large language models (LLMs). We demonstrate that the current generation of language models shows significant competence at deciphering cryptic crossword clues and outperforms previously reported state-of-the-art (SoTA) results by a factor of 2-3 in relevant benchmarks. We also develop a search algorithm that builds off this performance to tackle the problem of solving full crossword grids with out-of-the-box LLMs for the very first time, achieving an accuracy of 93% on New York Times crossword puzzles. Additionally, we demonstrate that LLMs generalize well and are capable of supporting answers with sound rationale.
@@
Soumadeep Saha, Sutanoya Chakraborty, Saptarshi Saha, Utpal Garain Indian Statistical Institute, Kolkata Correspondence: soumadeep.saha97@gmail.com
@@ 1 Introduction
 [^1] Code, data, etc., can be found at https://www.github.com/espressovi/LMCrossword
@@
Crossword puzzles are a type of word game that typically take the form of a square grid of white and black boxes. The objective of the puzzle is to fill the white boxes with letters from words or phrases based on the provided clues (see Figure 1). Although crosswords come in a variety of styles, the two most popular ones are the American style, or straight crosswords, and cryptic crosswords.
@@
Figure 1: Example of a crossword puzzle (left) and cryptic clues (right). (left) The grid must be filled up with answers from the semantic clues provided. The gray highlighted squares produce additional constraints, e.g., first character of the answer to clue 1 (across) and clue 4 (down) must be the same. Example by Fred Piscop. (right) In cryptic crosswords, the clues involve some form of wordplay and synonyms and often involve world knowledge. Examples are taken from the cryptonite dataset (see Appendix E for more).
Modern LLMs demonstrate astounding skills in reasoning, coding, wordplay, question answering (QA), and a multitude of other tasks (Wei et al., 2022a; Minaee et al., 2024). Despite the plethora of applications for LLMs seen today, their ability to generate language in a constrained setting remains relatively uncharted, and the ability to direct the generation process in order to meet certain criteria remains a challenge (Popescu-Belis et al., 2022). Consider the task of poem generation as an example, where, in addition to thematic aptness, constraints on rhyme or meter must be adhered to. Further, for certain kinds of poems like haikus, sonnets, or even song lyrics, restrictions on length, syllable counts, or patterns of stressed/unstressed syllables apply. Constraints also arise when dealing with formal languages (Koo et al., 2024), which is increasingly relevant in the LLM zeitgeist given their use in coding or interacting with databases, interpreters, APIs, etc. With a growing body of literature studying LLMs as agents (Madge and Poesio, 2024; Wang et al., 2024), language generation must follow environmental or physical constraints. For wider proliferation of LLM applications, they must demonstrate the ability to adhere to such domain-specific constraints (e.g., constraints imposed by knowledge graphs, tabular data, action spaces, etc.).
Solving crosswords requires proficiency in understanding contextual clues, semantics, wordplay, character manipulation, arithmetic, and reasoning (see Figure 1 (right)), along with satisfying constraints such as length limitations and character overlaps (see Figure 1). Given the multi-faceted nature of this task, it can serve as a testbed for studying constrained language generation. Since solving crosswords requires proficiency in several desirable areas, and identifying shortcomings can readily benefit other linguistic applications, in this paper we attempt to analyze LLMs’ ability to solve crossword puzzles, with the primary goal of understanding strengths and weaknesses demonstrated by SoTA LLMs. Our contributions in this study are as follows:
@@
$\diamond$ We perform extensive analysis to understand how well LLMs can answer crossword clues based on the provided contextual information and constraints. We show that current-generation SoTA LLMs demonstrate massively improved performance when compared to previous SoTA baselines, outperforming some previous benchmarks by a factor of 2 or 3, without any fine-tuning.
@@
$\diamond$ Our analysis shows that the ability of LLMs to adhere to length constraints, which is a critical component of solving crossword puzzles, is somewhat limited. We present results demonstrating that an LLM’s ability to count the number of constituent characters in a word degrades with the prevalence of the word. This suggests that their ability to perform this task does not generalize well. Additionally, we show that this is a significant limiting factor for crossword solving, since even when the generated answers are semantically similar, the length constraint is frequently not obeyed.
@@
$\diamond$ We devise a simple algorithm to tackle the full problem of solving crosswords with an LLM-guided search and incorporating grid information, solving straight crossword puzzles with an accuracy of 93.1%. This algorithm exploits the constraints imposed by previously generated answers to improve future answers and achieves results that are not far behind those of specialized automated crossword-solving systems. This algorithm also improves the clue-deciphering capability of LLMs, more than doubling the baseline performance.
$\diamond$ We perform human evaluation and further experiments with post knowledge-cutoff datasets to assess the soundness of logical reasoning, potential pitfalls, and generalizability. We find that SoTA LLMs appear to generalize well and provide sound logical reasoning with 74% of correct answers.
@@
We believe our work has been a major step towards demonstrating the power of LLMs with regard to crossword solving, and with future advances in certain key areas, LLMs’ performance in this task will be comparable to, if not better than, human experts and specialized systems. Our algorithmic approach of combining an LLM with a search strategy might be extended to linguistic problems with domain-specific constraints.
@@ 2 Background
Traditional approaches to solving straight crossword puzzles involve two key components: a candidate answer proposal system and a grid-filling algorithm (Angelini et al., 2005; Ginsberg, 2011). Answer proposal systems typically use similarity search on large clue-answer databases, fine-tuned language models, or a combination of both (Shazeer et al., 1999; Wallace et al., 2022). Grid-filling relies on variants of constraint satisfaction problem (CSP) algorithms. For instance, the system Proverb (Littman et al., 2002) achieves a 98.1% letter accuracy on New York Times (NYT) crosswords. Wallace et al. (2022) fine-tuned BERT and ByT5 on 6.4 million clue-answer pairs, and using a belief propagation algorithm, achieved a 99.7% letter accuracy. Kulshreshtha et al. (2022) established benchmarks with foundational language models and highlighted this task as “… a new high bar for AI systems. ”. Note that this work does not attempt to create an improved automated crossword-solving system but seeks to analyze foundational LLMs’ ability at this complex language task.
@@
Cryptic crosswords are more formidable and involve extensive wordplay such as anagrams, splicing, homophones, and puns. Traditional algorithms with large clue datasets and a context-free grammar parser (Deits, 2015) have shown poor performance, achieving only 7% accuracy (Sadallah et al., 2024). Recent studies have explored using large language models (LLMs) to solve cryptic crossword clues. Efrat et al. (2021) created a large dataset of cryptic crossword clues from UK dailies, The Times and The Telegraph, and fine-tuned a T5-Large (Raffel et al., 2020) model to establish baseline performances. They used a training split where answers in the training and test sets were mutually exclusive to prevent memorization. Rozner et al. (2021) curated a dataset from The Guardian and fine-tuned a T5-Large model using curriculum learning, showing performance improvements. They critiqued Efrat et al. (2021)’s approach, arguing that a disjoint train-test split is insufficient for teaching models to solve cryptic crosswords, as models exhibit “… robustness to plural and other inflections.” Instead, they proposed grouping similar-root words together in a split, noting that this more stringent criterion led to reduced performance.
@@
The most recent work by Sadallah et al. (2024) presents results with recent LLMs like Mistral-7B (Jiang et al., 2023), LLaMA2-7B (Touvron et al., 2023), and ChatGPT (OpenAI, 2021) in few-shot settings and also by fine-tuning the Mistral model. They report that ChatGPT surpasses other models with an accuracy of $9.5\%$, which demonstrates a significant performance gap with respect to human experts, who solve $99\%$ of cryptic crossword clues (Sadallah et al., 2024). They point out key limitations in their work, like the limited set of LLMs used and the potential for data contamination.
These recent works on cryptic crossword solving with LLMs highlight a significant performance gap between LLMs and human experts. However, these recent works approach the problem as a question answering (QA) task and ignore constraints imposed by the grid. This needs to be investigated further since it is yet to be seen whether a suitable approach that integrates constraint information into LLMs can yield significant performance benefits.
As for straight crosswords, Kulshreshtha et al. (2022) attempted to solve NYT crossword puzzles with language models and a Satisfiability Modulo Theory (SMT) solver with limited success, and resorted to culling the crossword grid based on candidate generations and ground truth answers. We present an algorithm employing LLM-guided search, which is significantly more successful at this task. With advancements in LLM capabilities and methods like ours, we believe LLMs could soon outperform humans in solving cryptic crosswords too.
@@ 3 Analyzing Crossword Clue Solving
The first step in solving a crossword puzzle is deciphering its clues, so following in the scheme of Rozner et al. (2021), Sadallah et al. (2024), Kulshreshtha et al. (2022), and Efrat et al. (2021), in this section, we first explore this as a QA task. We analyze the performance of several LLMs at different scales with variations of this task.
@@ 3.1 Datasets
Most of our analysis is performed on three crossword puzzle datasets. The first one covering straight or American-style crosswords was curated by us from the very popular and long-running puzzle section of the NYT. The other two, namely, Cryptonite by Efrat et al. (2021) and word-init-disjoint (abbreviated as Init) by Rozner et al. (2021), cover cryptic crossword puzzles. Note, the methodological differences between Cryptonite and Init, as discussed in Section 2, are not pertinent for us since we do not perform any training. A bulk of the results are reported on the NYT dataset we curated for straight crosswords and Init for cryptic crosswords since it was found to be more challenging than Cryptonite for LLMs. We drew 2000 randomly chosen samples to report results, and in-context examples were also randomly selected from a large pool of samples disjoint from the testing set. Further details can be found in Appendix B.
@@ 3.2 Answering Crossword Clues
Figure 2: Analyzing LLMs’ ability to generate answers from crossword clues. We test LLMs at different scales on the NYT, Cryptonite, and Init dataset with 5-shot, 10-shot prompts. All results are with T=0.5, further details about experimental protocols can be found in Appendix C.
In this experiment we only provided the LM with the query clue and length information (alongside instructions and in-context examples), with the expectation that it will produce the corresponding answer. We report results with Phi 3 3.8B Instruct (Abdin et al., 2024), Mistral 7B Instruct (Jiang et al., 2023), Llama 2 70B (Touvron et al., 2023), Llama 3 8B Instruct, Llama 3 70B (Meta, 2024), Mixtral 8x7B (Jiang et al., 2024), Claude 3 Sonnet (Anthropic, 2024), GPT 3.5 Turbo, and GPT 4 Turbo (OpenAI, 2023) to cover a wide range of parameter scales and a mix of open-weights and proprietary models. We investigate the performance of these models with few-shot prompts (5-shot, 10-shot) [^2] Further increases (25-shot) did not yield performance benefits. on samples from the NYT dataset (ours), Cryptonite, and the Init split (see Figure 2).
We find that the performance difference between 5-shot and 10-shot [^3] We elected not to perform experiments with 10-shot prompts on Claude and GPT-4-Turbo due to budget constraints. prompted answers is not significant (see Figure 2). We also note that there is an appreciable difference (up to 5%) in the performance of models across the board between Cryptonite and Init, with LLMs showing diminished performance on the Init dataset.
The models demonstrate improved performance with scale across datasets and, in particular, show remarkable improvement on the NYT dataset with Llama 3 70B, GPT 3.5 Turbo, Claude 3 Sonnet, and GPT-4-Turbo achieving 27.2%, 26.05%, 37.7% and 41.2% accuracy (exact match), respectively. The performance of LLMs on cryptic crosswords is worse compared to straight crosswords; however, Claude 3 Sonnet and GPT-4-Turbo outperform previous SoTA results on both the cryptic crossword datasets, achieving an accuracy of 12.9%, 23.5% on Cryptonite and 10.8%, 18.7% on Init respectively. The performance of GPT-4-Turbo is rather extraordinary, with a 1.97$\times$ improvement over previous SoTA (9.5% reported by Sadallah et al. (2024), see Figure 2).
@@ 3.3 Exploiting Partially Filled Grids
In the course of solving a crossword puzzle, we encounter intermediate states where some of the clues have been deciphered. Crossword solvers exploit these clues in order to inform their decisions about future answers. For example, in Figure 1 (left), when we want to solve for the clue in position 14 (across), we can use the characters from position 2 (down) and 4 (down) to narrow down the set of possible answers to only those that fit the template “_ T _ P”. In this section, we study LLMs’ performance at exploiting these constraints to arrive at better answers (see Table 1).
For this experiment we chose the best performing open-weights and proprietary models, i.e., LLaMA 3 70B and GPT-4-Turbo, respectively. We also report results on some smaller models to see if the trends hold. We report results on the NYT dataset and Init dataset with 5-shot prompts. For each query, $k$of the characters (letters) of the answer is provided alongside the clue and expected length of the answer, and the LLMs are expected to “unmask” the remaining characters using the provided constraint information and the crossword clue [^4] T=0.5, max-tokens=10, more details in Appendix C..
@@
Table 1: Testing if LLMs can improve by using character constraints from partially filled crosswords. Sadallah et al. (2024) reported an accuracy of 27.0% (70% hinted clues) by fine-tuning a Mistral 7B model on the Init dataset, which GPT-4-Turbo (76.30% accuracy) outperforms by a factor of $\sim$2.8$\times$ without fine-tuning.
We observe (Table 1) that, in all but one case, LLMs show improved performance with an increasing percentage of constraint information for both datasets. Additionally, to compare the performance of GPT-4-Turbo to previously reported SoTA results by Sadallah et al. (2024), we perform the experiment with the same settings and dataset split as them to find that GPT-4-Turbo with 5-shot prompts (76.3% accuracy) outperforms the fine-tuned Mistral 7B model (27% accuracy) by a factor of 2.8$\times$.
The fact that LLMs can successfully exploit constraints to answer crossword clues better suggests that they are well-suited to the task of solving full crosswords. The significant jump in performance observed in modern SoTA LLMs at this task and the task in Section 3.2 is extremely serendipitous, and we investigate this in further detail in Section 5.3.